OpenStudy (anonymous):

For breakfast, the cafeteria offers scrambled eggs with one slice of toast or an egg sandwich made with two slices of toast. In fifteen minutes, they sold 13 breakfasts and used 21 slices of bread. Use a model to determine how many egg sandwiches they sold. 6 7 8 9

OpenStudy (anonymous):

Simultaneous equations... let's say the number of scrambled eggs sold = a let's say the number of egg sandwiches sold = b a + b = total number of breakfasts sold = 13 so a+b=13 but we also know that every egg sandwich uses 2 slices. so total number of slices = a + 2(b) = 21 a + 2b = 21 can you solve from here?

OpenStudy (anonymous):

ya thanks!

OpenStudy (anonymous):

np :D